a) The origin in polar or cylindrical coordinates as compared to the rectangular coordinate system ______________________

A.

is fixed.

B.

none

C.

follows particle.

D.

is body centered.

b) If r = q 2 and q = 2t, find the magnitude of r and q when t = 2 seconds.

A.

4 cm/sec, 2 rad/sec2

B.

8 cm/sec, 16 rad/sec2

C.

16 cm/sec, 0 rad/sec2

D.

4 cm/sec, 0 rad/sec2

c) Cylindrical or polar coordinates are a suitable choice for the analysis of a problem for ____________________

A.

which data regarding the angular motion of the radial line r are given.

B.

none

C.

both A and B

D.

cases where the path can be conveniently expressed in terms of polar coordinates.
Answer #1

a)

Option A is correct.

The origin is fixed for polar as well as rectangular coordinates.

B)

Option C is correct .

Given, r = q^2 .

And q = 2t .

q'' = 0 , and t = 2 , q= 4 and r = 4^2 = 16 .

Therefore, option C is correct.

C)

Option A and D , ( Both are correct)

Cylindrical or polar coordinates are a suitable choice for the analysis of a problem forwhich data regarding the angular motion of the radial line r are given or cases where the path can be conveniently expressed in terms of polar coordinates.

If any doubt kindly comment. Thank you.